OpenStudy (anonymous):

The ratio of areas between two similar triangles is 1:4. if one side of the smaller triangle is 2 units, find the measure of the corresponding side of the other triangle.

hartnn (hartnn):

ratio of areas = (ratio of sides)^2

hartnn (hartnn):

so ratio of sides = ?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4} = (\frac{ 2 }{ x })^2\] what i got there is x=4 so the measure of the corresponding side of the trangle is 4 units?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thank you for the equation ! :))

hartnn (hartnn):

welcome ^_^